# Fake News in Sheep’s Clothing: Robust Fake News Detection Against LLM-Empowered Style Attacks

### v1

@@ Abstract.
It is commonly perceived that online fake news and reliable news exhibit stark differences in writing styles, such as the use of sensationalist versus objective language. However, we emphasize that style-related features can also be exploited for style-based attacks. Notably, the rise of powerful Large Language Models (LLMs) has enabled malicious users to mimic the style of trustworthy news outlets at minimal cost. Our analysis reveals that LLM-camouflaged fake news content leads to substantial performance degradation of state-of-the-art text-based detectors (up to 38% decrease in F1 Score), posing a significant challenge for automated detection in online ecosystems. To address this, we introduce SheepDog, a style-agnostic fake news detector robust to news writing styles. SheepDog achieves this adaptability through LLM-empowered news reframing, which customizes each article to match different writing styles using style-oriented reframing prompts. By employing style-agnostic training, SheepDog enhances its resilience to stylistic variations by maximizing prediction consistency across these diverse reframings. Furthermore, SheepDog extracts content-focused veracity attributions from LLMs, where the news content is evaluated against a set of fact-checking rationales. These attributions provide supplementary information and potential interpretability that assist veracity prediction. On three benchmark datasets, empirical results show that SheepDog consistently yields significant improvements over competitive baselines and enhances robustness against LLM-empowered style attacks.
@@
ccs: Information systems Data miningccs: Computing methodologies Natural language processing
@@ 1. Introduction
Automated fake news detectors are central to safeguarding integrity and harmony in the digital age (Nguyen et al., 2020; Shu et al., 2019; Wu and Hooi, 2023; Zhou et al., 2020). Style-related signals embedded in news articles serve as a crucial indicator of news veracity. For instance, reputable news sources uphold journalistic integrity, emphasize accuracy and fact-checking, and maintain a balanced tone (Bachmann et al., 2021). In stark contrast, unreliable outlets often employ sensationalism, lack credible sources, and may exhibit partisan biases (Higdon, 2020). Building upon this, prior advances have extracted sentiment features (Ajao et al., 2019; Zhang et al., 2021) to complement and enhance the detector, and highlighted the significance of writing style between hyperpartisan news and well-balanced mainstream reporting (Potthast et al., 2018).
@@
While style-related features are invaluable for discerning news veracity, they also offer a direct avenue for malicious users to conduct style-based attacks. This problem is exacerbated by the advent of powerful Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022), whose unprecedented capabilities for reasoning and generative tasks (Wei et al., 2022; Zheng and Zhan, 2023) reduces the gap between machine-generated and human-written news. Consequently, malicious actors now possess the capability to mimic the style of reputable news sources, in an attempt to evade automated detection. As shown in Figure 1, using a style-oriented prompt (i.e., “style of The New York Times”), LLM-camouflaged fake news successfully bypasses a RoBERTa (Liu et al., 2019)-based fake news detector.
@@
The impact of style-based attacks on fake news detectors has received limited attention, despite prior investigations into detector vulnerability related to attacks regarding social engagement (Chen and Shu, 2023), word-level perturbations (Koenders et al., 2021), and machine-generated malicious comments (Le et al., 2020). To assess the robustness of text-based fake news detectors, we introduce a series of style-based adversarial attacks, specifically by tailoring news articles to adversarial writing styles (details provided in Section 4.1). Our experiments, as summarized in Table 1, reveal a significant performance degradation in state-of-the-art text-based detectors, with some experiencing up to a 38% decline in F1 Score. These detectors, well-adapted to real news from trustworthy sources and fake news from unreliable sources, struggle to address the emerging real-world challenge where news content is presented in various styles.
@@
Motivated by the style-related vulnerability observed in state-of-the-art text-based detectors, we present SheepDog — a robust, style-agnostic fake news detector that effectively identifies deceptive content, even when concealed within the LLM-empowered ”sheep’s clothing”. Built upon a pretrained language model (LM) backbone, which can be fine-tuned to capture task-specific salient features, and leveraging the strong zero-shot reasoning and generative capabilities of LLMs, SheepDog capitalizes on the strengths of both (LLM: versatile; fine-tuned LM: specialized).
Our key ideas for overcoming style attacks are: (1) style-agnostic training, in which predictions across style-diverse news reframings are regulated to ensure maximal consistency, and (2) content-focused veracity attributions, in which textual insights on news veracity are extracted from an LLM to inform veracity predictions.
@@
The style-agnostic nature of SheepDog stems from its utilization of LLM-empowered news reframings. To encompass a wide spectrum of real-world news writing styles, we expand our training samples with style-diverse expressions, and ensure coherence across different expressions to establish style invariance. Specifically, through style-oriented prompts, we exploit the impressive text generation capabilities of LLMs to generate reframings for each news article, introducing a diversity of styles. Then, we maximize the consistency in the model’s veracity predictions across the original article and its various reframings. This strategy assists SheepDog to discount style-related features, which enables it to capture style-agnostic signals from the news content.
@@
To train a content-focused detector that is robust against stylistic variations, we further propose to distill the veracity-related knowledge from LLMs, specifically via eliciting content-focused attributions from inference APIs. We instruct the LLM to assess each article based on a defined set of content-focused fact-checking rationales (for comprehensive details, refer to Section 5.3). The resulting textual attributions serve to identify indicators of fake news. By converting these attributions into precise pseudo-labels, we introduce additional weak supervision that steers SheepDog towards learning a robust, style-agnostic representation of news. Our approach leverages these attribution-level predictions not only to facilitate style-agnostic training but also to potentially offer explainability into the veracity of news articles.
Overall, we make the following three contributions:
  • [leftmargin=*]
@@
  • LLM-Empowered Style Robustness: We introduce SheepDog, a style-agnostic fake news detector that achieves robustness through a combination of style-agnostic training and content-focused veracity attributions.
  • Effectiveness: Through extensive experiments, we demonstrate that leveraging LLM-empowered diverse news reframings significantly enhances the performance of representative language models (LMs) across three benchmark datasets.
@@ 2. Related Work
Fake News Detection. Automated fake news detection has been widely explored using a wide range of neural architectures (Devlin et al., 2019; Liu et al., 2019; Pelrine et al., 2021; Shu et al., 2019; Zhou et al., 2020). Apart from extracting lexical (Rashkin et al., 2017) and sentiment features (Potthast et al., 2018) within the news article text, many methods incorporate auxiliary features to supplement veracity prediction, including user comments (Shu et al., 2019), news environments (Sheng et al., 2022), knowledge bases (Cui et al., 2020; Dun et al., 2021), temporal patterns from users (Ruchansky et al., 2017), and social graphs (Lu and Li, 2020; Nguyen et al., 2020; Wu and Hooi, 2023). Recent studies also seek to address challenges including temporal shift (Hu et al., 2023), entity bias (Zhu et al., 2022a) and domain shift (Nan et al., 2021, 2022; Zhu et al., 2022b) in fake news detection scenarios. In this work, we adopt a text-based perspective and focus on news article texts, specifically on the stylistic variations in the news article under LLM-empowered style attacks.
@@
Adversarial Attack on Fake News Detectors. Investigating the vulnerabilities of fake news detectors is central to improving their real-world applicability. Hence, existing efforts (He et al., 2021a; Horne et al., 2019; Koenders et al., 2021; Lyu et al., 2023; Wang et al., 2023; Zhou et al., 2019) have studied the impact of different attacks from multiple aspects, including manipulation of social engagements (Lyu et al., 2023; Wang et al., 2023) and user behavior (He et al., 2021a), fact distortion (Koenders et al., 2021), subject-object exchange (Zhou et al., 2019), and blocking of data availability (Horne et al., 2019). However, the impact of writing styles remains underexplored. To bridge this gap, we investigate the robustness of text-based detectors against LLM-empowered style attacks, and propose a style-agnostic solution.
Large Language Models (LLMs). LLMs (OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023), with billions of parameters, have showcased remarkable reasoning capabilities, even matching or surpassing human performance in some scenarios (Wei et al., 2022; Zheng and Zhan, 2023). Benefiting from this success, recent work extracts knowledge from LLMs to facilitate a wide range of tasks, such as visual classification (Menon and Vondrick, 2023) and node classification (He et al., 2023). On a related front, LLM-generated text detection (Henrique et al., 2023; Muñoz-Ortiz et al., 2023) has also attracted increasing attention. Recent investigations further explore LLM-generated misinformation (Epstein et al., 2023; Huang et al., 2023; Zellers et al., 2019; Chen and Shu, 2023), and leverage LLMs for program-guided fact-checking (Pan et al., 2023). Although we also prompt an LLM to generate adversarial articles, we aim for style-agnostic detection that yields robust predictions across style-diverse news expressions. Moreover, existing work on LLM-generated misinformation mainly focuses on analysis of LLM-induced impacts and news article generation, which is orthogonal to our contributions. In this work, we focus on improving the detector’s style robustness by eliciting news reframings and content-focused veracity attributions from LLMs.
@@ 4.1. Attack Formulation
The impressive capabilities of LLMs (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022) have enabled malicious users to disguise fake news with restyling prompts, resulting in camouflaged articles that closely resemble reliable sources. Here, we explore a direct form of style-based attack utilizing news publisher names (e.g., ”The New York Times”). These names possess distinct styles that can be readily adopted by producers of fake news, making them a likely occurrence in real-world scenarios.
To replicate adversarial situations where news pieces are restyled in relation to various publishers, we manipulate the styles of both trustworthy and unreliable news. Specifically, in our test data, we utlize an LLM to rephrase real news in the style of tabloids, and fake news in the style of mainstream sources. Our general prompt format is shown as follows: {tcolorbox}[colback=black!2!white,colframe=white!50!black,boxrule=0.5mm] Rewrite the following article using the style of [publisher name]: [news article] Based on publisher popularity, in the place of [publisher name], we select “National Enquirer” to camouflage real news, and “The New York Times” to camouflage fake news. These LLM-camouflaged test articles are then employed to evaluate the resilience of a detector against style-based attacks, as illustrated in Figure 1.
@@ 4.2. Style-Related Detector Vulnerability
Automated fake news detection becomes increasingly difficult with LLM-empowered style attacks. In this subsection, we conduct preliminary analysis on real-world news to evaluate the influence of writing styles on text-based detectors. Our analysis is based on the FakeNewsNet (Shu et al., 2020) benchmark (consisting of PolitiFact and GossipCop datasets) and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017). Further details regarding these benchmark datasets can be found in Table 8 and Section 6.1.1. Specifically, we seek to address the following question: To what extent can text-based fake news detectors withstand LLM-empowered style attacks?
In Table 1, we examine $13$ representative text-based detectors under both original and adversarial settings (detailed method descriptions are provided in Section 6.1.2). These detectors employ diverse model architectures, including Convolutional Neural Networks (CNNs) (Zhou et al., 2020), Recurrent Neural Networks (RNNs), (Shu et al., 2019), Graph Neural Networks (GNNs) applied to document graphs (Vaibhav et al., 2019), Transformer-based LMs (Devlin et al., 2019; He et al., 2021b; Liu et al., 2019; Zhang et al., 2021; Schick and Schütze, 2021; Hu et al., 2022a; Xie et al., 2020), and LLMs (OpenAI, 2022; Ouyang et al., 2022; Touvron et al., 2023). We evaluate model robustness against LLM-empowered style attack based on performance on the adversarial test set formulated in Section 4.1. Our empirical results yield the following insights:
@@
Observation 2 (Insufficiency of LLMs as Zero-Shot Fake News Detectors).
@@
Our two findings reveal a critical challenge in achieving robustness again stylistic variations. Ideally, an effective fake news detector should prioritize content assessment over style. While the news content typically remains stable, style can undergo rapid changes in the dynamic online environment. Our observations, however, reveals a stark deviation from this ideal. Detectors overly influenced by style may struggle to reliably differentiate between real and fake news, and even powerful LLMs may prove inadequate for the specific demands of fake news detection. The accessibility for malicious users to manipulate style using LLMs further exacerbates this issue, emphasizing the pressing need for a style-agnostic approach.
Table 1. Under LLM-empowered style attacks, existing text-based fake news detectors suffer severe performance deterioration in terms of F1 Score (%). Performance degradation is computed on adversarial test set (O: original; A: adversarial).
@@
Figure 2. Overview of our SheepDog style-agnostic fake news detector.

### v2

@@ Abstract.
It is commonly perceived that fake news and real news exhibit distinct writing styles, such as the use of sensationalist versus objective language. However, we emphasize that style-related features can also be exploited for style-based attacks. Notably, the advent of powerful Large Language Models (LLMs) has empowered malicious actors to mimic the style of trustworthy news sources, doing so swiftly, cost-effectively, and at scale. Our analysis reveals that LLM-camouflaged fake news content significantly undermines the effectiveness of state-of-the-art text-based detectors (up to 38% decrease in F1 Score), implying a severe vulnerability to stylistic variations. To address this, we introduce SheepDog, a style-robust fake news detector that prioritizes content over style in determining news veracity. SheepDog achieves this resilience through (1) LLM-empowered news reframings that inject style diversity into the training process by customizing articles to match different styles; (2) a style-agnostic training scheme that ensures consistent veracity predictions across style-diverse reframings; and (3) content-focused veracity attributions that distill content-centric guidelines from LLMs for debunking fake news, offering supplementary cues and potential intepretability that assist veracity prediction. Extensive experiments on three real-world benchmarks demonstrate SheepDog’s style robustness and adaptability to various backbones.  [^1] Data and code are available at: https://github.com/jiayingwu19/SheepDog.
@@
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671977isbn: 979-8-4007-0490-1/24/08ccs: Information systems Data miningccs: Computing methodologies Natural language processing
@@ 1. Introduction
Psychological theories, such as the Undeutsch hypothesis (Amado et al., 2015), suggest that genuine and fake statements exhibit distinct linguistic styles. Indeed, reputable news sources uphold journalistic integrity, emphasize accuracy and fact-checking, and maintain a balanced tone (Bachmann et al., 2021). In contrast, unreliable outlets often resort to sensationalism, lack credible sources, and may exhibit partisan biases (Higdon, 2020). Building upon these stylistic differences, recent advances in automated fake news detection have incorporated sentiment features (Ajao et al., 2019; Zhang et al., 2021) to enhance the detector, and highlighted the significance of styles in discerning between hyperpartisan news and well-balanced mainstream reporting (Potthast et al., 2018).
@@
While style-related features serve as key indicators in identifying fake news, they also offer a direct avenue for malicious users to conduct style-based attacks. This problem is exacerbated by the advent of powerful Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022), whose unprecedented capabilities for reasoning and generative tasks (Wei et al., 2022; Zheng and Zhan, 2023) bridges the gap between machine-generated and human-written news. Consequently, malicious actors now possess the capability to mimic the style of reputable news sources, in an attempt to evade automated detection. As shown in Figure 1, using a style-oriented prompt (i.e., “style of The New York Times”), LLM-camouflaged fake news successfully bypasses a RoBERTa (Liu et al., 2019)-based fake news detector.
@@
The impact of stylistic variations on fake news detectors has received limited attention, despite prior investigations into attacks regarding social engagements (Wang et al., 2023), word-level perturbations (Koenders et al., 2021), and machine-generated malicious comments (Le et al., 2020). To assess the robustness of text-based fake news detectors, we introduce a series of style-based attacks, specifically by tailoring news articles to adversarial writing styles (detailed in Section 4.1). Our experiments, as summarized in Table 1, reveal a significant performance degradation in state-of-the-art text-based detectors, with some suffering up to a 38% decline in F1 Score. Existing detectors, well-fitted to real news from trustworthy sources and fake news from unreliable sources, struggle to adapt to real-world scenarios where news content is presented in diverse styles.
@@
To address the style-related vulnerability of text-based detectors, we introduce SheepDog, a style-robust approach that consistently recognizes trustworthy content and identifies deceptive content, even when concealed within the LLM-empowered “sheep’s clothing”. Built upon a pretrained language model (LM) backbone, which can be fully fine-tuned to capture task-specific salient features, and leveraging the strong zero-shot reasoning and generative capabilities of LLMs, SheepDog effectively combines the strengths of both (fine-tuned LM: specialized; LLM: versatile).
@@
The style-agnostic nature of SheepDog stems from its utilization of LLM-empowered news reframings. To accommodate the wide spectrum of writing styles presented in real-world news articles, we harness the impressive capabilities of LLMs in adhering to complex real-world instructions (He et al., 2024). By doing so, we can reframe training articles into style-diverse expressions while preserving the integrity of their content. Subsequently, through style-agnostic training, we aim to ensure consistency in the model’s veracity predictions across each news article and its style-diverse reframings. This training scheme encourages SheepDog to discount style-related features, enabling it to focus on capturing style-agnostic veracity signals from the news content.
@@
To reinforce the emphasis of our approach on news content over style, we propose to incorporate content-focused veracity attributions from an LLM to inform veracity predictions. Leveraging the extensive world knowledge and reasoning capabilities within LLMs (Asai et al., 2024; Menon and Vondrick, 2023), we elicit explanatory outputs from LLMs regarding the veracity of news articles, with reference to a set of content-centric fake news debunking rationales specifically related to news content (detailed in Section 5.3). By converting these rationales into precise pseudo-labels, we introduce additional weak supervision that steers SheepDog towards learning robust, style-agnostic news representations. Our approach leverages these attribution-level predictions not only to facilitate style robustness but also to potentially offer explainability into the veracity of news articles.
Our key contributions are summarized as follows:
@@
  • LLM-Empowered Style Robustness: We introduce SheepDog, a style-agnostic fake news detector that achieves robustness through style-agnostic training and content-focused veracity attribution prediction, synergized within a multi-task learning paradigm.
  • Effectiveness: Extensive experiments demonstrate that SheepDog achieves significantly superior style robustness across multiple style-based adversarial settings, and yields consistent performance gains when combined with representative LM and LLM backbones.
@@ 2. Related Work
Fake News Detection. Automated fake news detection has been explored using a wide range of neural architectures (Pelrine et al., 2021; Shu et al., 2019; Zhou et al., 2020). Apart from extracting lexical (Rashkin et al., 2017) and sentiment features (Potthast et al., 2018) within the news article text, many methods incorporate auxiliary features to supplement veracity prediction, including user comments (Shu et al., 2019), news environments (Sheng et al., 2022), knowledge bases (Cui et al., 2020; Dun et al., 2021), temporal patterns from users (Ruchansky et al., 2017), and social graphs (Nguyen et al., 2020; Wu and Hooi, 2023; Wu et al., 2023). Recent studies also seek to address challenges including temporal shift (Hu et al., 2023), entity bias (Zhu et al., 2022a) and domain shift (Nan et al., 2021, 2022; Zhu et al., 2022b) in fake news detection scenarios. In this work, we adopt a text-based perspective, specifically focusing on enhancing the robustness of fake news detectors against stylistic variations.
@@
Adversarial Attack on Fake News Detectors. Investigating the vulnerabilities of fake news detectors is central to improving their real-world applicability. Hence, existing efforts (He et al., 2021a; Horne et al., 2019; Koenders et al., 2021; Lyu et al., 2023; Wang et al., 2023; Zhou et al., 2019) have studied the impact of different attacks from multiple aspects, including manipulation of social engagements (Lyu et al., 2023; Wang et al., 2023) and user behavior (He et al., 2021a), fact distortion (Koenders et al., 2021), subject-object exchange (Zhou et al., 2019), and blocking of data availability (Horne et al., 2019). However, the impact of writing styles remains underexplored. To bridge this gap, we investigate the resilience of text-based detectors against LLM-empowered style attacks, and propose a style-agnostic solution.
LLM Capabilities and Misinformation. LLMs (OpenAI, 2022; Touvron et al., 2023; OpenAI, 2023) have demonstrated remarkable reasoning capabilities that even match or surpass human performance in certain scenarios (Wei et al., 2022; Zheng and Zhan, 2023). However, the impressive strengths of LLMs have also attracted increasing attention towards LLM-generated misinformation (Kreps et al., 2022). Recent investigations have found that LLMs can act as high-quality misinformation generators (Huang et al., 2023; Lucas et al., 2023; Pan et al., 2023a; Zellers et al., 2019), and that LLM-generated misinformation is generally harder to detect (Chen and Shu, 2023, 2024). On a related front, recent work explore the role of LLMs as fact-checkers (Guan et al., 2023; Pan et al., 2023b; Zhang and Gao, 2023) and fake news detectors (Chen and Shu, 2024; Pelrine et al., 2023), and leverage the commonsense reasoning capabilities to elicit supplementary explanations from LLMs (Asai et al., 2024; He et al., 2023; Hu et al., 2024; Menon and Vondrick, 2023) that facilitate a wide range of tasks. Although we also instruct an LLM to generate style-related adversarial articles, our goal is to simulate real-world scenarios where news are presented in diverse styles. Additionally, instead of leveraging LLMs to make veracity judgments that distinguish false information from the truth (Chen and Shu, 2024; Hu et al., 2024; Lucas et al., 2023; Pelrine et al., 2023), in this work, we investigate the role of LLMs in enhancing the style robustness of text-based fake news detectors, specifically through injecting style-diverse reframings and content-centric cues into the training process.
@@ 4.1. Attack Formulation
The impressive capabilities of LLMs (Brown et al., 2020; OpenAI, 2022; Ouyang et al., 2022) enable malicious users to disguise fake news with restyling prompts, resulting in camouflaged articles that closely resemble reliable sources. In this work, we explore a direct form of style-based attack utilizing news publisher names (e.g., “CNN”). These names possess distinct styles that can be readily adopted by producers of fake news, making them a likely occurrence in real-world scenarios.
To simulate the adversarial situations where news articles are restyled in relation to various publishers, we manipulate the styles of both trustworthy and unreliable news. Specifically, among the test samples, we utlize an LLM to rephrase real news in the style of tabloids, and fake news in the style of mainstream sources. Our general prompt format is shown as follows:
Rewrite the following article using the style of [publisher name]: [news article]
Based on publisher popularity, in the place of [publisher name], we select “National Enquirer” to transform real news, and “CNN” to transform fake news. These LLM-restyled test articles are then employed to evaluate the resilience of a detector against style-based attacks, as illustrated in Figure 1.
@@ 4.2. Style-Related Detector Vulnerability
Automated fake news detection becomes increasingly difficult against LLM-empowered style attacks. In this subsection, we conduct preliminary analysis on real-world new articles to evaluate the influence of writing styles on text-based detectors. Our analysis is based on the FakeNewsNet (Shu et al., 2020) benchmark (consisting of PolitiFact and GossipCop datasets) and the Labeled Unreliable News (LUN) dataset (Rashkin et al., 2017), with dataset descriptions relegated to Section 6.1.1 and Table 2. Specifically, we investigate the following question: To what extent can text-based fake news detectors withstand LLM-empowered style attacks?
In Table 1, we examine $13$ representative text-based detectors under both original and adversarial settings (detailed method descriptions are provided in Section 6.1.2). These detectors encompass three categories: (1) text-based fake news detectors with diverse task-specific architectures, including Recurrent Neural Networks (RNNs) (Shu et al., 2019), Convolutional Neural Networks (CNNs) (Zhou et al., 2020), Graph Neural Networks (GNNs) applied to document graphs (Vaibhav et al., 2019), and Transformers (Zhang et al., 2021); (2) Fine-tuned LMs (Devlin et al., 2019; He et al., 2021b; Liu et al., 2019; Schick and Schütze, 2021; Hu et al., 2022a; Xie et al., 2020) on the fake news detection benchmark datasets; and (3) LLMs (OpenAI, 2022; Ouyang et al., 2022; Touvron et al., 2023) with zero-shot prompting. Few-shot and fine-tuned LLM experiments are relegated to Appendix A.
We evaluate the robustness of detectors against LLM-empowered style attack based on their performance under the adversarial setting outlined in Section 4.1. Our empirical results in Table 1 and Appendix A yield the following two implications:
@@
Observation 2 (Insufficiency of LLMs as Fake News Detectors).
@@
Our two findings suggest a fundamental limitation of text-based fake news detectors in achieving robust veracity predictions against stylistic variations. Detectors overly influenced by styles struggle to reliably differentiate between real and fake news, and even the powerful LLMs may prove inadequate for the specific demands of fake news detection. In the dynamic digital landscape, the styles of news articles evolve rapidly, while the accessibility for malicious users to manipulate style using LLMs exacerbates these variations. Therefore, for effective deployment, a fake news detector must prioritize the assessment of news content over style. This objective motivates our subsequent innovations toward a style-agnostic fake news detection approach.
Table 1. Under LLM-empowered style attacks, existing text-based fake news detectors suffer severe performance deterioration in terms of F1 Score (%). (O: original; A ($\downarrow$): gap between original unperturbed performance and adversarial performance on the test set formulated in Section 4.1).
@@
Figure 2. Overview of the proposed SheepDog framework for style-agnostic fake news detection.